RC Meeting and Run Plan, Monday, February 10th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

Monday 10th Feb. 2020.

  • Production at 3.4 GeV until about 7am. At that point the intrinsic beam size was too wide (Ion Chamber reading too high). Since there was no time to retune the beam before the 8:30 access, we finished the last 1h30 taking Carbon optics data. Then at 8:30, polarized beam was terminated and we changed target cells from Dutch → Big Brother.

General updates

Pending problem

Recent polarization measurements

Current Run Plan

  • Monday 00:00 2:30am: continue ⊥ production
  • 2:30am NMR, then switch to //
  • 2:30am → 8:30am do // production
  • 8:30: Beam is terminated. Ask MCC to send RadCon for full Hall survey
  • 8:30 → whenever the Hall survey is finished: NMR systematic studies:
    • Go to pick-up coil position
    • Do NMR (pol //, lasers on)
    • Turn off lasers. Wait 15min for target temperature to stabilize.
    • Do NMR (pol //, lasers off)
    • Turn on lasers wait for 15min for temperature to stabilize.
    • Do NMR (//, laser on).
    • Turn off lasers. Wait 15min for target temperature to stabilize.
    • Do NMR (pol //, lasers off)
    • Rotate to ⊥ then do NMR (pol ⊥, lasers off)
    • Rotate to // then do NMR (pol //, lasers off)
    • Rotate to ⊥ then do NMR (pol ⊥, lasers off)
    • Turn on lasers. Wait 15min for target temperature to stabilize.
    • Do NMR (pol ⊥, lasers on)
    • Repeat the laser off sequence if time allows. Otherwise, start Dutch → Big Brother target change procedure.
  • Plan for the Dutch → Big Brother target change

  • If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change the beam tuning (except work on transport magnets), the do beam recovery procedure (do YAG viewer part only if your are suspicious of the harp scan results)
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM

!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!

  • General idea: take 6 hours transverse data to get polarization up, then 10 hours longitudinal data


  • Replace target cell at some point (Monday Feb. 10th ?)
  • Replacement of Kepco PS to be done at the same time.
  • Replace the 2 old flowmeters for the nitrogen cooling jets lines with 2 new ones (Andy K.)
  • In the same access: Replace the EPR fiber bundle (Jian-Ping)

Hall Access

When there is a down:

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If SHMS dipole trips again, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]
  • Do not change SHMS momentum until we fixed the current lead valve

Planned Accelerator Configuration Changes

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links